Two equal charges are separated 75cm apart in free air. Calculate the quantity of either charge if the force between them is 0.17N. What will be the quantity if the charges are situated in an insulating medium of dielectric constant 10?

1
Expert's answer
2022-02-04T06:37:24-0500

1.

"q=r\sqrt \frac Fk=3.26~\mu C,"

2.

"q'=q\sqrt\varepsilon=10.31~\mu C."